# Chevalley formulae for the motivic Chern classes of Schubert cells and for the stable envelopes

### v3

@@ 7.2. Changing the polarizations
A natural question is to study the change of the stable envelopes when we vary the above three parameters. To start, the change of chambers is encoded in the left Weyl group action. More precisely, the group $G$ acts on $T^{*}(G/B)$ by left multiplication, which induces a left Weyl group action on $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$, see [MNS22b]. If we change the chamber $\mathfrak{C}$ to another chamber $w(\mathfrak{C})$ ($w\in W$), we have the following formula (see [AMSS19, Lemma 8.2(a)])
@@ 7.3. Changing the alcoves
The alcoves in $\mathfrak{t}^{*}_{\mathbb{R}}$ are of the form $x(\operatorname{A_{\circ}})+\lambda$ for some $x\in W$ and some $\lambda$ in the root lattice. It was proved in [AMSS19, Lemma 8.2],[SZZ20, Rmk. 2.3] that
@@
Let $Z=T^{*}(G/B)\times_{\mathcal{N}}T^{*}(G/B)$ be the Steinberg variety where ${\mathcal{N}}\subset\mathfrak{g}$ denotes the nilpotent cone. There is an algebra isomorphism due to Kazhdan–Lusztig and Ginzburg [KL87, CG09]
@@
[AMSS19, Theorem 8.6] Let $\iota:G/B\hookrightarrow T^{*}(G/B)$ denote the inclusion of the zero section. For any $w\in W$,
@@ 8.1. Whittaker functions
In this section we study Whittaker functions. These appear in $p$-adic representation theory, and in this note we utilize a cohomological construction of these functions from [AMSS19], see also [MSA22]. Recall the definition of the Demazure–Lusztig operators on $K_{T}(pt)[y]$:
@@
The operators satisfy the usual quadratic and braid relations in the Hecke algebra, therefore for any $w\in W$ there are operators $\widetilde{T_{w}}$ and $\widetilde{T^{\vee}_{w}}$ acting on $K_{T}(pt)[y]$, and defined using any reduced decomposition of $w$. The following has been proved in [MSA22, Thm. 1.1]:
@@
where ${\mathcal{W}}_{\lambda,w}$ is the Iwahori–Whittaker function for the Langlands dual group over a non-archimedean local field; we refer to [MSA22] for more details, including the number theoretic definition of ${\mathcal{W}}_{\lambda,w}$. Here we identified $y$ with $-q^{-1}$, where $q$ is the number of elements in the residue field. As explained in [MSA22], from the fact that
@@
Here $\chi_{w_{0}\lambda}$ denotes the character for the irreducible representation of $G$ of highest weight $w_{0}(\lambda)$. We also note that, in type A, an interpretation of the Iwahori–Whittaker function in terms of the partition function of the Iwahori lattice model has been obtained in [BBBG19].
@@
where the second equality follows from [AMSS19, Theorem 4.2(3)]. Therefore, taking the equivariant Euler characteristics of both sides of Equation 27, we get
@@
As a corollary we prove a variant of the Casselman–Shalika formula (38), obtained by Li [Li92], see also [BBBG19, Proposition 9.4]. First define
@@ 8.2. Hall–Littlewood polynomials
Equation 43 follows because $\lambda_{y}(T^{*}_{G/P})=MC_{y}(G/P)=\displaystyle\sum_{w\in W^{P}}MC_{y}(X(wW_{P})^{\circ})$, Theorem 5.15, and $\chi_{T}(MC_{y}(X(uW_{P})^{\circ}))=(-y)^{\ell(u)}$ for any $u\in W^{P}$. Equation 44 follows from the localization formula (cf. [Nie74], [MSA22, Theorem 2.1 (c)].) To be more specific, the torus fixed points in $G/P_{\lambda}$ are $\{wP_{\lambda}\mid w\in W^{P}\}$, and the torus weights of the tangent space at the fixed point $wP_{\lambda}$ are $\{-w\alpha\mid\alpha\in R^{+}\setminus R^{+}_{\lambda}\}$. ∎
@@ Appendix A Chevalley formulae for the Chern–Schwartz–MacPherson classes of Schubert cells
In this appendix, we give a short proof of the Chevalley formulae for the equivariant Chern–Schwartz–MacPherson (CSM) and Segre–MacPherson (SM) classes of the Schubert cells in the (partial) flag varieties, see [AMSS17, Thm  9.10] and Theorem A.4 below. Our proof again relies on the action of the appropriate Hecke algebra, this time on the equivariant cohomology of $G/P$. Besides the intrinsic interest in these Chevalley formulae, we mention that recursions based on it were recently utilized to obtain proofs of Nakada’s colored hook formula for finite Weyl groups [MNS22a] and more general Coxeter groups [MNS24].
@@ A.2. Definition of the CSM/SM classes
Next we recall the basic definitions and properties of CSM classes for the Schubert cells in $G/P$, we will be brief. We refer the reader e.g. to [Ohm06, AMSS17] for details, including a construction of these classes in the equivariant setting and for general varieties.
[…]
@@ A.3. The Chevalley formula in cohomology
We now specialize to $X=G/P$ with the usual $T$-action. For simplicity we will denote by $s_{\mathrm{M}}(Z\subset G/P)$ simply by $s_{\mathrm{M}}(Z)$. We will identify the equivariant (Borel-Moore) homology group $H_{*}^{T}(G/P)$ with the equivariant cohomology $H^{*}_{T}(G/P)$, using the Poincaré duality. The sets of CSM classes of Schubert cells $\{c_{\mathrm{SM}}(X(wW_{P})^{\circ})\mid w\in W^{P}\}$ and of the SM classes $\{s_{\mathrm{M}}(X(wW_{P})^{\circ})\mid w\in W^{P}\}$ form bases for $H_{T}^{*}(G/P)_{\operatorname{loc}}:=H_{T}^{*}(G/P)\otimes_{H_{T}^{*}(pt)}\operatorname{Frac}H_{T}^{*}(pt)$. Moreover, if one takes the opposite Schubert cells in any of these sets, then the two bases are dual under the usual intersection pairing, see [AMSS17, Theorem 9.4]:
@@
Recall for any $\lambda\in X^{*}(T)_{P}$, ${\mathcal{L}}_{\lambda}:=G\times^{P}{\mathbb{C}}_{\lambda}\in\operatorname{Pic}_{T}(G/P)$. The following is our main result in this Appendix, and it has also been proved in [AMSS17, Thm. 9.10] using the Chevalley formula for the cohomological stable envelopes from [Su16, Thm. 3.7]. Here we give a direct proof based on the action of the degenerate affine Hecke algebra.
@@ Appendix B An example of the $\lambda$-chain formula
(a) Alcove walk $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}+\lambda$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}p_{w}=c^{-}_{2}c^{-}_{1}$ , $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\lambda}=c^{+}_{0}c^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{+}_{0}$
@@
(a) Alcove walk $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}+\lambda$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}p_{w}=c^{-}_{2}c^{-}_{1}$ , $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{\lambda}=c^{+}_{0}c^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{+}_{0}$
(b) Alcove walk $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{-\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}-\lambda$$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}p_{w}=c^{-}_{2}c^{-}_{1}$, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}p_{-\lambda}=c^{-}_{2}c^{-}_{1}c^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}$
@@
$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr p_{w}\tilde{p}_{\lambda}&\tilde{p}_{\lambda}&\mathcal{M}&J&\varphi(p)=u&{\rm wt}(p)=w\tilde{r}^{\lambda}_{J_{>}}(\lambda)\\ \hline\cr\hline\cr(0)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}c^{-}_{1}c^{-}_{0}c^{-}_{2}c^{+}_{0}}&\{\}&\{\}&s_{2}s_{1}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(1)&f^{+}_{0}c^{-}_{2}c^{+}_{1}c^{-}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1}\}&\{6\}&s_{1}&-\varpi_{1}+\varpi_{2}\\ (2)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}}f^{+}_{1}c^{-}_{0}c^{+}_{2}c^{+}_{0}&\{h_{3}\}&\{4\}&s_{1}&-\varpi_{2}\\ (3)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}c^{-}_{1}c^{-}_{0}}f^{+}_{2}c^{+}_{0}&\{h_{5}\}&\{2\}&s_{1}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(4)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{-}_{0}&\{h_{2}\}&\{5\}&s_{2}&2\varpi_{1}-2\varpi_{2}\\ (5)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}c^{-}_{1}}f^{+}_{0}c^{+}_{2}c^{-}_{0}&\{h_{4}\}&\{3\}&s_{2}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(6)&f^{+}_{0}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1},h_{2}\}&\{5,6\}&id&\varpi_{1}\\ (7)&f^{+}_{0}c^{-}_{2}c^{+}_{1}f^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1},h_{4}\}&\{3,6\}&id&-\varpi_{1}+\varpi_{2}\\ (8)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}f^{+}_{0}&\{h_{2},h_{6}\}&\{1,5\}&id&2\varpi_{1}-2\varpi_{2}\\ (9)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}}f^{+}_{1}f^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{3},h_{4}\}&\{3,4\}&id&-\varpi_{2}\\ (10)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}c^{-}_{0}c^{-}_{2}c^{-}_{1}}f^{+}_{0}c^{+}_{2}f^{+}_{0}&\{h_{4},h_{6}\}&\{1,3\}&id&\varpi_{1}-3\varpi_{2}\\ \hline\cr\end{array}$

### v4

@@ 7.2. Changing the polarizations
A natural question is to study the change of the stable envelopes when we vary the above three parameters. To start, the change of chambers is encoded in the left Weyl group action. More precisely, the group $G$ acts on $T^{*}(G/B)$ by left multiplication, which induces a left Weyl group action on $K_{T\times{\mathbb{C}}^{*}}(T^{*}(G/B))$, see [MNS22b]. If we change the chamber $\mathfrak{C}$ to another chamber $w(\mathfrak{C})$ ($w\in W$), we have the following formula (see [AMSS24a, Lemma 8.2(a)])
@@ 7.3. Changing the alcoves
The alcoves in $\mathfrak{t}^{*}_{\mathbb{R}}$ are of the form $x(\operatorname{A_{\circ}})+\lambda$ for some $x\in W$ and some $\lambda$ in the root lattice. It was proved in [AMSS24a, Lemma 8.2],[SZZ20, Rmk. 2.3] that
@@
Let $Z=T^{*}(G/B)\times_{\mathcal{N}}T^{*}(G/B)$ be the Steinberg variety where ${\mathcal{N}}\subset\mathfrak{g}$ denotes the nilpotent cone. There is an algebra isomorphism due to Kazhdan–Lusztig and Ginzburg [KL87, CG10]
@@
[AMSS24a, Theorem 8.6] Let $\iota:G/B\hookrightarrow T^{*}(G/B)$ denote the inclusion of the zero section. For any $w\in W$,
@@ 8.1. Whittaker functions
In this section we study Whittaker functions. These appear in $p$-adic representation theory, and in this note we utilize a cohomological construction of these functions from [AMSS24a], see also [MS22]. Recall the definition of the Demazure–Lusztig operators on $K_{T}(pt)[y]$:
@@
The operators satisfy the usual quadratic and braid relations in the Hecke algebra, therefore for any $w\in W$ there are operators $\widetilde{T_{w}}$ and $\widetilde{T^{\vee}_{w}}$ acting on $K_{T}(pt)[y]$, and defined using any reduced decomposition of $w$. The following has been proved in [MS22, Thm. 1.1]:
@@
where ${\mathcal{W}}_{\lambda,w}$ is the Iwahori–Whittaker function for the Langlands dual group over a non-archimedean local field; we refer to [MS22] for more details, including the number theoretic definition of ${\mathcal{W}}_{\lambda,w}$. Here we identified $y$ with $-q^{-1}$, where $q$ is the number of elements in the residue field. As explained in [MS22], from the fact that
@@
Here $\chi_{w_{0}\lambda}$ denotes the character for the irreducible representation of $G$ of highest weight $w_{0}(\lambda)$. We also note that, in type A, an interpretation of the Iwahori–Whittaker function in terms of the partition function of the Iwahori lattice model has been obtained in [BBBG24].
@@
where the second equality follows from [AMSS24a, Theorem 4.2(3)]. Therefore, taking the equivariant Euler characteristics of both sides of Equation 27, we get
@@
As a corollary we prove a variant of the Casselman–Shalika formula (36), obtained by Li [Li92], see also [BBBG24, Proposition 9.4]. First define
@@ 8.2. Hall–Littlewood polynomials
Equation 41 follows because $\lambda_{y}(T^{*}_{G/P})=MC_{y}(G/P)=\displaystyle\sum_{w\in W^{P}}MC_{y}(X(wW_{P})^{\circ})$, Theorem 5.15, and $\chi_{T}(MC_{y}(X(uW_{P})^{\circ}))=(-y)^{\ell(u)}$ for any $u\in W^{P}$. Equation 42 follows from the localization formula (cf. [Nie74], [MS22, Theorem 2.1 (c)].) To be more specific, the torus fixed points in $G/P_{\lambda}$ are $\{wP_{\lambda}\mid w\in W^{P}\}$, and the torus weights of the tangent space at the fixed point $wP_{\lambda}$ are $\{-w\alpha\mid\alpha\in R^{+}\setminus R^{+}_{\lambda}\}$. ∎
@@ Appendix A Chevalley formulae for the Chern–Schwartz–MacPherson classes of Schubert cells
In this appendix, we give a short proof of the Chevalley formulae for the equivariant Chern–Schwartz–MacPherson (CSM) and Segre–MacPherson (SM) classes of the Schubert cells in the (partial) flag varieties, see [AMSS23, Thm  9.10] and Theorem A.4 below. Our proof again relies on the action of the appropriate Hecke algebra, this time on the equivariant cohomology of $G/P$. Besides the intrinsic interest in these Chevalley formulae, we mention that recursions based on it were recently utilized to obtain proofs of Nakada’s colored hook formula for finite Weyl groups [MNS22a] and more general Coxeter groups [MNS24].
@@ A.2. Definition of the CSM/SM classes
Next we recall the basic definitions and properties of CSM classes for the Schubert cells in $G/P$, we will be brief. We refer the reader e.g. to [Ohm06, AMSS23] for details, including a construction of these classes in the equivariant setting and for general varieties.
The (additive) group of constructible functions ${\mathcal{F}}(X)$ consists of functions $\varphi=\sum_{Z}c_{Z}1\hskip-3.5pt1_{Z}$, where the sum is over a finite set of constructible subsets $W\subset X$, $c_{Z}\in{\mathbb{Z}}$ are integers, and $1\hskip-3.5pt1_{Z}$ is the characteristic function of $Z$. For a proper morphism $f:Y\rightarrow X$, there is a linear map $f_{*}:{\mathcal{F}}(Y)\rightarrow{\mathcal{F}}(X)$, such that for any constructible subset $Z\subset Y$, $f_{*}(1\hskip-3.5pt1_{Z})(x)=\chi_{\textit{top}}(f^{-1}(x)\cap Z)$, where $x\in X$ and $\chi_{\textit{top}}$ denotes the topological Euler characteristic. A conjecture attributed to Deligne and Grothendieck states that there is a unique natural transformation $c_{*}:{\mathcal{F}}\to H_{*}$ from the functor of constructible functions on a complex algebraic variety $X$ to the homology functor, where all morphisms are proper, such that if $X$ is smooth then $c_{*}(1\hskip-3.5pt1_{X})=c(T_{X})\cap[X]$. This conjecture was proved by MacPherson [Mac74]; the class $c_{*}(1\hskip-3.5pt1_{X})$ for possibly singular $X$ was shown to coincide with a class defined earlier by M.-H. Schwartz [Sch65a, Sch65b, BS81].
@@ A.3. The Chevalley formula in cohomology
We now specialize to $X=G/P$ with the usual $T$-action. For simplicity we will denote by $s_{\mathrm{M}}(Z\subset G/P)$ simply by $s_{\mathrm{M}}(Z)$. We will identify the equivariant (Borel-Moore) homology group $H_{*}^{T}(G/P)$ with the equivariant cohomology $H^{*}_{T}(G/P)$, using the Poincaré duality. The sets of CSM classes of Schubert cells $\{c_{\mathrm{SM}}(X(wW_{P})^{\circ})\mid w\in W^{P}\}$ and of the SM classes $\{s_{\mathrm{M}}(X(wW_{P})^{\circ})\mid w\in W^{P}\}$ form bases for $H_{T}^{*}(G/P)_{\operatorname{loc}}:=H_{T}^{*}(G/P)\otimes_{H_{T}^{*}(pt)}\operatorname{Frac}H_{T}^{*}(pt)$. Moreover, if one takes the opposite Schubert cells in any of these sets, then the two bases are dual under the usual intersection pairing, see [AMSS23, Theorem 9.4]:
@@
Recall for any $\lambda\in X^{*}(T)_{P}$, ${\mathcal{L}}_{\lambda}:=G\times^{P}{\mathbb{C}}_{\lambda}\in\operatorname{Pic}_{T}(G/P)$. The following is our main result in this Appendix, and it has also been proved in [AMSS23, Thm. 9.10] using the Chevalley formula for the cohomological stable envelopes from [Su16, Thm. 3.7]. Here we give a direct proof based on the action of the degenerate affine Hecke algebra.
@@ Appendix B An example of the $\lambda$-chain formula
(a) Alcove walk $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}p_{\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}+\lambda$
$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p_{w}=c^{-}_{2}c^{-}_{1}$ , $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}p_{\lambda}=c^{+}_{0}c^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{+}_{0}$
@@
(b) Alcove walk $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}p_{-\lambda}$ from $\operatorname{A_{\circ}}$ to $\operatorname{A_{\circ}}-\lambda$
$\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}p_{w}=c^{-}_{2}c^{-}_{1}$, $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}p_{-\lambda}=c^{-}_{2}c^{-}_{1}c^{-}_{2}c^{-}_{0}c^{-}_{1}c^{-}_{2}$
@@
$\begin{array}[]{|c|c|c|c|c|c|}\hline\cr p_{w}\tilde{p}_{\lambda}&\tilde{p}_{\lambda}&\mathcal{M}&J&\varphi(p)=u&{\rm wt}(p)=w\tilde{r}^{\lambda}_{J_{>}}(\lambda)\\ \hline\cr\hline\cr(0)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}c^{-}_{2}c^{-}_{1}c^{-}_{0}c^{-}_{2}c^{+}_{0}}&\{\}&\{\}&s_{2}s_{1}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(1)&f^{+}_{0}c^{-}_{2}c^{+}_{1}c^{-}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1}\}&\{6\}&s_{1}&-\varpi_{1}+\varpi_{2}\\ (2)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}c^{-}_{2}}f^{+}_{1}c^{-}_{0}c^{+}_{2}c^{+}_{0}&\{h_{3}\}&\{4\}&s_{1}&-\varpi_{2}\\ (3)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}c^{-}_{2}c^{-}_{1}c^{-}_{0}}f^{+}_{2}c^{+}_{0}&\{h_{5}\}&\{2\}&s_{1}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(4)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{-}_{0}&\{h_{2}\}&\{5\}&s_{2}&2\varpi_{1}-2\varpi_{2}\\ (5)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}c^{-}_{2}c^{-}_{1}}f^{+}_{0}c^{+}_{2}c^{-}_{0}&\{h_{4}\}&\{3\}&s_{2}&\varpi_{1}-3\varpi_{2}\\ \hline\cr(6)&f^{+}_{0}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1},h_{2}\}&\{5,6\}&id&\varpi_{1}\\ (7)&f^{+}_{0}c^{-}_{2}c^{+}_{1}f^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{1},h_{4}\}&\{3,6\}&id&-\varpi_{1}+\varpi_{2}\\ (8)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}}f^{+}_{2}c^{+}_{1}c^{+}_{0}c^{+}_{2}f^{+}_{0}&\{h_{2},h_{6}\}&\{1,5\}&id&2\varpi_{1}-2\varpi_{2}\\ (9)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}c^{-}_{2}}f^{+}_{1}f^{+}_{0}c^{+}_{2}c^{+}_{0}&\{h_{3},h_{4}\}&\{3,4\}&id&-\varpi_{2}\\ (10)&{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}c^{-}_{0}c^{-}_{2}c^{-}_{1}}f^{+}_{0}c^{+}_{2}f^{+}_{0}&\{h_{4},h_{6}\}&\{1,3\}&id&\varpi_{1}-3\varpi_{2}\\ \hline\cr\end{array}$